HGH-RESOLUTION NEAR-INFRARED ABSORPTIONS of CH_{4}/He PLASMAS

\underline{MAN-CHOR CHAN}, SHUN-HING YEUNG, YUNGFENG Li, YUEN-YEE WONG, KAN-HING YIM, and WANG-MING CHAN , Department of Chemistry, The Chinese University of Hong Kong, Shatin, N. T., Hong Kong.

Near infrared absorptions ( 11000-13000 cm^{−1}) of a gaseous plasma generated by discharging a gas mixture of He and CH_{4} (at a ratio of 220:1 to give a total pressure of 860 mTorr) was studied using a high-resolution Ti:sapphire laser spectrometer. Using a home-built AC high voltage power supply and a 2 m long multi-traversal hollow cathode discharge cell equilibrated at 210 K, a peak-to-peak discharge current of 1-1.5 A was maintained for the plasma. The absorption signals were detected using phase-sensitive detection with concentration modulation at 5-10 kHz. In addition to the absorptions due to the Phillips band system (^{1}P i_{u} – ^{1}Sigma_{g} electronic transitions) of C_{2},^{a} several unidentified bands were also observed with good signal-to-noise ratio. Based on the earlier work of Ballik and Ramsey,^{b} and S. P. Davis et al.,^{c} the analysis of the spectrum is underway.

aJ. G. Phillips, Ap. J. 107, 389 (1948).

bJ. G. Phillips, Ap. J. 137, 84 (1963).

cS. P. Davis, M. C. Abrams, J. G. Phillips, and M. L. P. Rao, J. Opt. Soc. Am. B5, 2280 (1988).